OpenStudy (anonymous):

What is the slope of the line between the points (–3, 5) and (5, 1)?

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (anonymous):

slopes btwn points (x1,y1) and (x2,y2) is given by (y2-y1)/(x2-x1)

hartnn (hartnn):

@caught , so what did u get m=... ?

OpenStudy (anonymous):

-2 dont know if that correct

hartnn (hartnn):

only slope was asked..

hartnn (hartnn):

no, not -2, but close...

OpenStudy (hba):

@hartnn My bad lol :P

hartnn (hartnn):

what u got y1-y2 =.... ?

OpenStudy (hba):

m=y_2-y_1/x_2-x_1 (x_1,y_1)=(–3, 5) and (x_2,y_2)= (5, 1)?

OpenStudy (precal):